OpenStudy (anonymous):

If F(theta)=tan theta=3, find F(theta+pi)

OpenStudy (lgbasallote):

hint: \[F(\theta + \pi) = \tan(\theta + \pi)\] does that help?

OpenStudy (anonymous):

not really

OpenStudy (lgbasallote):

okay...you have \[\tan \theta = 3\] do you know how to solve for theta?

OpenStudy (anonymous):

no thats what i need to know

OpenStudy (lgbasallote):

take arctangent of 3..what do you get?

OpenStudy (anonymous):

-0.143

OpenStudy (anonymous):

is that wrong

OpenStudy (anonymous):

please help me my teacher didnt go over this part

OpenStudy (anonymous):

How did -0.143 come about?

OpenStudy (anonymous):

i put it in my calculator

OpenStudy (anonymous):

now im really confused

OpenStudy (anonymous):

\[\tan ^{-1} (3) = 1.249\] ..?